# A HHO formulation for variable density incompressible flows where the density is purely advected

### v1

@@ 3.4 Semi-discrete HHO formulation
$+\int_{\partial T^{\rm D}}\bigl((\bm{u}_{\partial T}-\bm{g}_{\rm D})\otimes\bm{n}_{\partial T}\bigr):\left(\nu{\mathcal{G}}_{T}\underline{\bm{v}}_{T}+\frac{\nu}{h_{F}}\bm{v}_{\partial T}\otimes\bm{n}_{\partial T}\right)+$
$-\int_{\partial T^{\rm D}}\nu{\mathcal{G}}_{T}\underline{\bm{u}}_{T}:(\bm{v}_{\partial T}\otimes\bm{n}_{\partial T}),$ (8a)
$sd^{\rm sms}_{T}$ $\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{z}_{T}\bigr)\coloneqq sd^{\rm sms}_{T,{\rm!bc}}(\underline{\bm{u}}_{T},\underline{\rho}_{T};\underline{z}_{T})$
@@
$+\int_{\partial T^{\rm N}}\bigl(\bm{u}_{\partial T}\cdot\bm{n}_{\partial T}\bigr)^{\oplus}\bigl(\rho_{\partial T}{z}_{\partial T}\bigr)+\int_{\partial T^{\rm N}}\bigl(\bm{u}_{\partial T}\cdot\bm{n}_{\partial T}\bigr)^{\ominus}\bigl(g_{\rm In}{z}_{\partial T}\bigr)$
@@
$+\int_{\partial T^{\rm D}}\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\oplus}\bigl(\rho_{\partial T}{z}_{\partial T}\bigr)+\int_{\partial T^{\rm D}}\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\ominus}\bigl(g_{\rm In}{z}_{\partial T}\bigr),$ (8b)
$sd^{{\rm vol}}_{T}$ $(\underline{\bm{u}}_{T};\underline{q}_{T})\coloneqq sd^{{\rm vol}}_{T,{\rm!bc}}(\underline{\bm{u}}_{T};\underline{q}_{T})-\int_{\partial T^{\rm N}}(\bm{u}_{\partial T}\cdot\bm{n}_{\partial T})\,{q}_{\partial T}-\int_{\partial T^{\rm D}}(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T})\,{q}_{\partial T}.$ (8c)
We remark that $\bm{u}_{\partial T}\cdot\bm{n}_{\partial T}=\bm{u}_{T}\cdot\bm{n}_{\partial T}$ on $\partial T_{\rm N}$, see Section 3.5 for details.
@@
$sd_{T}^{{\rm mnt}}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr)\coloneqq\int_{T}\frac{\partial({\rho}_{T}\bm{u}_{T})}{\partial t}\cdot\bm{v}_{T}+sd^{{\rm slm}}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr),$
$sd_{T}^{\rm mss}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{z}_{T}\bigr)\coloneqq\int_{T}\frac{\partial\rho_{T}}{\partial t}z_{T}+h_{T}\int_{\partial T}\left(\frac{\partial\rho_{\partial T}}{\partial t}-\frac{\partial{\rho}_{T}}{\partial t}\right)({z}_{\partial T}-{z}_{T})+sd^{\rm sms}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{z}_{T}\bigr).$ (9)
We remark that the second term of (9) is a time derivative stabilization designed to guarantee the robustness of the formulation whenever the normal velocity component happens to be null with respect to a mesh face. In this event, thanks to the additional stabilizations, the face density definition falls back to the average of the densities of the two elements sharing the face. We remark that a similar strategy was proposed to stabilize the linear momentum equations in the limit of vanishing viscosity, see [7] and [4].
The global residuals of the spatial discretization $r^{\rm mnt}_{h}((\underline{\bm{u}}_{h},\underline{\rho}_{h},\underline{p}_{h});\cdot):\underline{\bm{V}}^{k}_{h}\to\mathbb{R}$, $r^{\rm mss}_{h}((\underline{\bm{u}}_{h},\underline{\rho}_{h});\cdot):\underline{Z}^{k}_{h}\to\mathbb{R}$, and $r^{\rm vol}_{h}(\underline{\bm{u}}_{h};\cdot):\underline{Q}^{k+1}_{h}\to\mathbb{R}$ are obtained by cell-by-cell assembly of the local discretizations, i.e.: For all $\underline{\bm{v}}_{h}\in\underline{\bm{V}}_{h}^{k}$, all $\underline{z}_{h}\in\underline{Z}_{h}^{k}$ and all $\underline{q}_{h}\in\underline{Q}_{h}^{k+1}$,
$r_{h}^{\rm mnt}\bigl((\underline{\bm{u}}_{h},\underline{\rho}_{h},\underline{p}_{h});\underline{\bm{v}}_{h}\bigr)$ $\coloneqq\sum_{T\in\mathcal{T}_{h}}sd^{\rm mnt}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr),$ (10a)
@@
$r_{h}^{\rm mnt}\bigl((\underline{\bm{u}}_{h},\underline{\rho}_{h},\underline{p}_{h});\underline{\bm{v}}_{h}\bigr)$ $=0$ $\forall\underline{\bm{v}}_{h}\in\underline{\bm{V}}_{h}^{k},$ (11)
@@ 3.5 H(div)-conformity and independence from irrotational body forces
Following [8] we consider the HHO formulation of the volume conservation equation. The requirement $r_{h}^{\rm vol}(\underline{\bm{u}}_{h};\underline{q}_{h})=0$ for each $\underline{\bm{u}}_{h}\in\underline{\bm{V}}_{h}$ and for all $\underline{q}_{h}\in\underline{Q}^{k+1}_{h}$ reads
$0=\sum_{T\in\mathcal{T}_{h}}\left(-\int_{T}{q}_{T}\;\nabla\cdot\bm{u}_{T}+\int_{\partial T}{q}_{\partial T}\bm{u}_{T}\cdot\bm{n}_{TF}-\int_{\partial T^{\rm N}}{q}_{\partial T}\bm{u}_{\partial T}\cdot\bm{n}_{\partial T}-\int_{\partial T^{\rm D}}{q}_{\partial T}\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\right).$
@@
    $\forall F\in\mathcal{F}_{h}^{\rm D},$ $\int_{F}{q}_{F}(\bm{u}_{T}-\bm{g}_{\rm D}){\cdot}\bm{n}_{TF}=0,$ $\;\text{i.e., }\;\bm{u}_{T}{\cdot}\bm{n}_{TF}{=}\pi_{F}^{k+1}(\bm{g}_{D}{\cdot}\bm{n}_{TF}),$
@@
    where the equalities involving normal velocity components hold true $\forall\bm{x}\in F$. Note that a Neumann face is such that $\bm{u}_{F}\in\mathcal{P}^{k+1}(F)$. Note that an internal face is such that $F=\partial T\cap\partial T^{\prime}$, with $^{\prime}\in\mathcal{T}_{h}, TTT\neq T^{\prime}$ and $\bm{n}_{TF}=-\bm{n}_{T^{\prime}F}$.
The above procedure allows to conclude that the velocity field is pointwise divergence-free and $\bm{H}(\operatorname{div})$-conforming, and to infer that conservation of volume is enforced cell-by-cell.
@@
$\int_{T}\bm{f}\cdot\bm{v}_{T}$ $=\int_{T}\bm{\ell}\cdot\bm{v}_{T}+\int_{T}(\nabla\psi)\cdot\bm{v}_{T}$ (12)
$=\int_{T}\bm{\ell}\cdot\bm{v}_{T}-\int_{T}\psi\;(\nabla\cdot\bm{v}_{T})+\int_{\partial T}\psi\;\bm{v}_{T}\cdot\bm{n}_{\partial T}$
$=\int_{T}\bm{\ell}\cdot\bm{v}_{T}-\int_{T}\pi^{k}_{T}\psi\;(\nabla\cdot\bm{v}_{T})+\int_{\partial T}\pi^{k+1}_{\partial T}\psi\;\bm{v}_{T}\cdot\bm{n}_{\partial T}$
$=\int_{T}\bm{\ell}\cdot\bm{v}_{T}+\int_{T}{\bm{\mathfrak{g}}}_{T}^{k+1}(\underline{I}_{{Q},T}\psi)\cdot\bm{v}_{T},\quad$
where, for all $q\in H^{1}(\Omega)$, $\underline{I}_{\underline{Q},T}q\coloneqq\bigl(\pi_{T}^{k}q,\pi_{\partial T}^{k+1}q\bigr)$, is the local interpolation operator onto $\underline{Q}_{T}$. Note that integration by parts allows to pass from the first to the second line. Moreover, since $\bm{v}_{T}\in\mathcal{P}^{k+1}(T)$, $\nabla\cdot\bm{v}_{T}\in\mathcal{P}^{k}(T)$, and $\bm{n}_{TF}$ is a constant vector over $F$, $L^{2}$-projectors can be introduced on the third line.
Introducing (12) in $sd^{{\rm slm}}_{T,{\rm!bc}}$, see in particular (5), it becomes clear that the irrotational part of the body force can be combined with the pressure gradient discretization to obtain an equivalent reformulation of the momentum equation spatial dicretization. The procedure required to achieve the desired reformulation can be summarized as follows: replace
@@
    $\int_{T}{\bm{\mathfrak{g}}}_{T}^{k+1}\underline{p}_{T}\cdot\bm{v}_{T}$ with $\int_{T}{\bm{\mathfrak{g}}}_{T}^{k+1}(\underline{p}_{T}-\underline{I}_{{Q},T}\psi)\cdot\bm{v}_{T}$
@@
    $\int_{T}\bm{f}\cdot\bm{v}_{T}$ with $\int_{T}\bm{\ell}\cdot\bm{v}_{T}$,
@@
    -$\int_{\partial T^{\rm N}}p_{\partial T}\bm{v}_{\partial T}\cdot\bm{n}_{\partial T}$ with $\int_{\partial T^{\rm N}}(p_{\partial T}-\pi_{\partial T}^{k+1}\psi)\bm{v}_{\partial T}\cdot\bm{n}_{\partial T}$
@@
    $\int_{\partial T^{\rm N}}\bm{g}_{\rm N}\cdot\bm{v}_{\partial T}$ with $\int_{\partial T^{\rm N}}(\bm{g}_{\rm N}-\pi_{\partial T}^{k+1}\psi\bm{n}_{\partial T})\cdot\bm{v}_{\partial T}$

### v3

@@ 3.4 Semi-discrete HHO formulation
$+\int_{\partial T^{\rm D}}\bigl((\bm{u}_{\partial T}-\bm{g}_{\rm D})\otimes\bm{n}_{\partial T}\bigr):\left(\nu{\mathcal{G}}_{T}\underline{\bm{v}}_{T}+\frac{\nu}{h_{F}}\bm{v}_{\partial T}\otimes\bm{n}_{\partial T}\right)$
$-\int_{\partial T^{\rm D}}\nu{\mathcal{G}}_{T}\underline{\bm{u}}_{T}:(\bm{v}_{\partial T}\otimes\bm{n}_{\partial T})-\int_{\partial T^{\rm N}}p_{\partial T}(\bm{v}_{\partial T}\cdot\bm{n}_{\partial T})+\int_{\partial T^{\rm N}}\bm{g}_{\rm N}\cdot\bm{v}_{\partial T},$ (11a)
$sd^{\rm sms}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{z}_{T}\bigr)$ $\coloneqq sd^{\rm sms}_{T,{\rm!bc}}(\underline{\bm{u}}_{T},\underline{\rho}_{T};\underline{z}_{T})-\frac{1}{2}\int_{\partial T^{{\rm D},{\rm N}}}\bigl(\bm{u}_{T}\cdot\bm{n}_{\partial T}\bigr)\bigl(\rho_{\partial T}{z}_{\partial T}\bigr)$
@@
$+\int_{\partial T^{\rm N}}\left(\bigl(\bm{u}_{\partial T}\cdot\bm{n}_{\partial T}\bigr)^{\oplus}\rho_{\partial T}+\bigl(\bm{u}_{\partial T}\cdot\bm{n}_{\partial T}\bigr)^{\ominus}g_{\rm In}\right){z}_{\partial T}$
@@
$+\int_{\partial T^{\rm D}}\left(\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\oplus}\rho_{\partial T}+\bigl(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\bigr)^{\ominus}g_{\rm In}\right){z}_{\partial T},$ (11b)
$sd^{{\rm vol}}_{T}(\underline{\bm{u}}_{T};\underline{q}_{T})$ $\coloneqq sd^{{\rm vol}}_{T,{\rm!bc}}(\underline{\bm{u}}_{T};\underline{q}_{T})-\int_{\partial T^{\rm N}}(\bm{u}_{\partial T}\cdot\bm{n}_{\partial T})\,{q}_{\partial T}-\int_{\partial T^{\rm D}}(\bm{g}_{\rm D}\cdot\bm{n}_{\partial T})\,{q}_{\partial T},$
where $\alpha^{\oplus}\coloneqq\frac{1}{2}(\alpha+|\alpha|)$ and $\alpha^{\ominus}\coloneqq\frac{1}{2}(\alpha-|\alpha|)$.
@@
$sd_{T}^{{\rm mnt},\pi^{p}}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr)\coloneqq\int_{T}\left({\rho}_{T}\frac{\partial\bm{u}_{T}}{\partial t}+\frac{\bm{u}_{T}}{2}\frac{\partial{\rho}_{T}}{\partial t}\right)\cdot\bm{v}_{T}+sd^{{\rm slm},\pi^{p}}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr),$ (12)
$sd_{T}^{\rm mss}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{z}_{T}\bigr)\coloneqq\int_{T}\frac{\partial\rho_{T}}{\partial t}z_{T}+\frac{h_{T}}{(k+1)^{2}}\int_{\partial T}\left(\frac{\partial{\rho}_{T}}{\partial t}-\frac{\partial\rho_{\partial T}}{\partial t}\right)({z}_{T}-{z}_{\partial T})+sd^{\rm sms}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T});\underline{z}_{T}\bigr).$ (13)
We remark that the time derivative in the momentum equation has been modified in view of achieving stability, as first proposed by [17], see Sec. 3.7 for details. We further remark that the second term of (13) is a time derivative stabilization designed to guarantee the robustness of the formulation whenever the normal velocity component vanishes on a mesh face, see Sec. 3.6 for additional details.
The global residuals of the spatial discretization $r^{{\rm mnt},\pi^{p}}_{h}((\underline{\bm{u}}_{h},\underline{\rho}_{h},\underline{p}_{h});\cdot):\underline{\bm{V}}^{k}_{h}\to\mathbb{R}$, $r^{\rm mss}_{h}((\underline{\bm{u}}_{h},\underline{\rho}_{h});\cdot):\underline{Z}^{k}_{h}\to\mathbb{R}$, and $r^{\rm vol}_{h}(\underline{\bm{u}}_{h};\cdot):\underline{Q}^{k+1}_{h}\to\mathbb{R}$ are obtained by cell-by-cell assembly of the local discretizations, i.e.: For all $\underline{\bm{v}}_{h}\in\underline{\bm{V}}_{h}^{k}$, all $\underline{z}_{h}\in\underline{Z}_{h}^{k}$ and all $\underline{q}_{h}\in\underline{Q}_{h}^{k+1}$,
$r_{h}^{{\rm mnt},\pi^{p}}\bigl((\underline{\bm{u}}_{h},\underline{\rho}_{h},\underline{p}_{h});\underline{\bm{v}}_{h}\bigr)$ $\coloneqq\sum_{T\in\mathcal{T}_{h}}sd^{{\rm mnt},\pi^{p}}_{T}\bigl((\underline{\bm{u}}_{T},\underline{\rho}_{T},\underline{p}_{T});\underline{\bm{v}}_{T}\bigr),$ (14a)
@@
$r_{h}^{{\rm mnt},\pi^{p}}\bigl((\underline{\bm{u}}_{h},\underline{\rho}_{h},\underline{p}_{h});\underline{\bm{v}}_{h}\bigr)$ $=0$ $\forall\underline{\bm{v}}_{h}\in\underline{\bm{V}}_{h}^{k},$ (15)
@@ 3.5 H(div)-conformity and independence from irrotational body forces
Let’s focus on the HHO discretization of the volume conservation equation. The requirement $r_{h}^{\rm vol}(\underline{\bm{u}}_{h};\underline{q}_{h})=0$ for each $\underline{\bm{u}}_{h}\in\underline{\bm{V}}_{h}$ and for all $\underline{q}_{h}\in\underline{Q}^{k+1}_{h}$ reads
$0=-\sum_{T\in\mathcal{T}_{h}}\int_{T}{q}_{T}\;\nabla\cdot\bm{u}_{T}+\sum_{T\in\mathcal{T}_{h}}\left(\int_{\partial T}{q}_{\partial T}\bm{u}_{T}\cdot\bm{n}_{TF}-\int_{\partial T^{\rm N}}{q}_{\partial T}\bm{u}_{\partial T}\cdot\bm{n}_{\partial T}-\int_{\partial T^{\rm D}}{q}_{\partial T}\bm{g}_{\rm D}\cdot\bm{n}_{\partial T}\right).$
@@
    $\forall F\in\mathcal{F}_{h}^{\rm D},$ $\int_{F}{q}_{F}(\bm{u}_{T}-\bm{g}_{\rm D}){\cdot}\bm{n}_{TF}=0,$ $\;\text{i.e., }\;\bm{u}_{T}{\cdot}\bm{n}_{TF}{=}\pi_{F}^{k+1}(\bm{g}_{D}{\cdot}\bm{n}_{TF}),$ (16)
@@
    where the equalities involving the normal velocity component hold true $\forall\bm{x}\in F$. Note that a Neumann face is such that $\bm{u}_{F}\in\mathcal{P}^{k+1}(F)$. Note that an internal face is such that $F=\partial T\cap\partial T^{\prime}$, with $^{\prime}\in\mathcal{T}_{h}, TTT\neq T^{\prime}$ and $\bm{n}_{TF}=-\bm{n}_{T^{\prime}F}$.
The above constraints ensure that the velocity field is pointwise divergence-free and $\bm{H}(\operatorname{div})$-conforming, and allow to infer that conservation of volume is enforced cell-by-cell.
@@
$\int_{T}\bm{f}\cdot\bm{v}_{T}$ $=\int_{T}\bm{\ell}\cdot\bm{v}_{T}+\int_{T}(\nabla\psi)\cdot\bm{v}_{T}=\int_{T}\bm{\ell}\cdot\bm{v}_{T}-\int_{T}\psi\;(\nabla\cdot\bm{v}_{T})+\int_{\partial T}\psi\;\bm{v}_{T}\cdot\bm{n}_{\partial T}$ (17)
$=\int_{T}\bm{\ell}\cdot\bm{v}_{T}-\int_{T}\pi^{k}_{T}\psi\;(\nabla\cdot\bm{v}_{T})+\int_{\partial T}\pi^{k+1}_{\partial T}\psi\;\bm{v}_{T}\cdot\bm{n}_{\partial T}=\int_{T}\bm{\ell}\cdot\bm{v}_{T}+\int_{T}{\bm{\mathfrak{g}}}_{T}^{k+1}(\underline{I}_{{Q},T}\psi)\cdot\bm{v}_{T},\quad$
where, for all $q\in H^{1}(\Omega)$, $\underline{I}_{\underline{Q},T}q\coloneqq\bigl(\pi_{T}^{k}q,\pi_{\partial T}^{k+1}q\bigr)$, is the local interpolation operator onto $\underline{Q}_{T}$. Note that integration is employed on the last step of the first line. Moreover, since $\bm{v}_{T}\in\mathcal{P}^{k+1}(T)$, $\nabla\cdot\bm{v}_{T}\in\mathcal{P}^{k}(T)$, and $\bm{n}_{TF}$ is a constant vector over $F$, $L^{2}$-projectors can be introduced on the second line.
Introducing (17) in $sd^{{\rm slm},\pi^{p}}_{T,{\rm!bc}}$, see in particular (6), it becomes clear that the irrotational part of the body force can be combined with the pressure gradient discretization to obtain an equivalent reformulation of the momentum equation spatial dicretization. The procedure required to achieve the desired reformulation can be summarized as follows: replace
@@
    $\int_{T}{\bm{\mathfrak{g}}}_{T}^{k+1}\underline{p}_{T}\cdot\bm{v}_{T}$ with $\int_{T}{\bm{\mathfrak{g}}}_{T}^{k+1}(\underline{p}_{T}-\underline{I}_{{Q},T}\psi)\cdot\bm{v}_{T}$ in (6),
@@
    $\int_{T}\bm{f}\cdot\bm{v}_{T}$ with $\int_{T}\bm{\ell}\cdot\bm{v}_{T}$ in (6),
@@
    -$\int_{\partial T^{\rm N}}p_{\partial T}\bm{v}_{\partial T}\cdot\bm{n}_{\partial T}$ with $\int_{\partial T^{\rm N}}(p_{\partial T}-\pi_{\partial T}^{k+1}\psi)\bm{v}_{\partial T}\cdot\bm{n}_{\partial T}$ in (11a),
@@
    $\int_{\partial T^{\rm N}}\bm{g}_{\rm N}\cdot\bm{v}_{\partial T}$ with $\int_{\partial T^{\rm N}}(\bm{g}_{\rm N}-\pi_{\partial T}^{k+1}\psi\bm{n}_{\partial T})\cdot\bm{v}_{\partial T}$ in (11a).